# Generative User-Experience Research for Developing Domain-specific Natural Language Processing Applications

### v1

@@ Abstract
User experience (UX) is a part of human-computer interaction (HCI) research and focuses on increasing intuitiveness, transparency, simplicity, and trust for the system users. Most of the UX research for machine learning (ML) or natural language processing (NLP) focuses on a data-driven methodology, i.e., it fails to focus on users’ requirements, and engages domain users mainly for usability evaluation. Moreover, more typical UX methods tailor the systems towards user usability unlike learning about the user needs first. The paper proposes a methodology for integrating generative UX research into developing domain NLP applications. Generative UX research employs domain users at the initial stages of prototype development, i.e., ideation and concept evaluation, and the last stage for evaluating the change in user value. In the case study, we report the full-cycle prototype development of a domain-specific semantic search for daily operations in the process industry. Our case study shows that involving domain experts increases their interest and trust in the final NLP application. Moreover, we show that synergetic UX+NLP research efficiently considers data- and user-driven opportunities and constraints, which can be crucial for NLP applications in narrow domains.
@@
generative research, natural language processing, human-computer interaction, user experience, domain-specific language, semantic search
@@ 1 Introduction
Often the data-driven approach falls short in exploring the needs of the domain users [51]. On the one hand, domain users are often integrated into development at the late test phase to evaluate the usability of ML/NLP applications [13]. Unlike user-driven software development, the development of NLP applications depends mainly on data availability or experimenting with machine learning (ML)/NLP trends and thus is a major driver of application development. On the other hand, the user-driven development of a domain-specific ML/NLP application in medicine showed that close collaboration with the domain users in the earlier stages increases the effectiveness of the final product [50]. Therefore, integrating user experience (UX) and human-computer interaction (HCI) research into ML/NLP research addresses users’ needs, fuses their expertise, and increases intuitiveness, transparency, simplicity, and trust for the system users [9, 37].
@@
This paper proposes a methodology for integrating generative research into a development cycle of domain NLP applications by combining data- and user-driven concept creation, prototype development, and prototype evaluation. Widely applied in design and software development, generative UX research is a methodology that aims to learn about domain users and their stories [27, 5]. To increase credibility and value in the final application, generative research engages the users in the early stage of development. Prototyping or developing the first minimal viable products (MVPs) most benefits from the methodology. The proposed methodology consists of the following phases. First, the methodology describes generative UX research steps to explore opportunities for ideas and application concepts. The proposed generative research for NLP applications relies on a combination of exploratory data analysis and contextual inquiries, i.e., user interviews and on-site observations, as a source of idea generation. Furthermore, the domain users validate a proposed concept and help define the most important features for prototype implementation. Second, the prototyping addresses the most important identified research challenges, which are crucial for usability assessment. Lastly, the prototype evaluation measures user value via the accuracy evaluation and repeated contextual inquiries. Therefore, the methodology combines data- and user-driven approaches to full-cycle UX-driven prototype development.
The paper reports a case study of applying the proposed methodology. We develop a real-world domain NLP application for improving effectiveness and efficiency in daily operations in the process industry. We apply generative research combined with exploratory data analysis to identify where an NLP application could help the best based on the use cases and stories of domain users. Our NLP prototype is integrated into an industrial logging system, which tracks daily operations and events in a plant. The goal is to increase users’ interest and trust in the application. The paper reports a detailed case study, which could be reused as a step-by-step guideline for applied NLP research.
@@
Our team comprises UX researchers, data scientists, software developers, and a project leader. The team composition facilitated a deeper understanding of the capabilities and limitations of the domain NLP application [18]. Therefore, the paper may have specific interest for the individual practitioners with the above-mentioned roles or similar teams. Moreover, practitioners of user-driven research, e.g., customer research departments, and quality assurance officers could benefit from the proposed methodology and reported results.
@@
The paper is structured as follows: Section 3 proposes a full-cycle prototype development via generative research, from generative research for concept ideation to the user-driven prototype evaluation. Section 4 reports a methodology and results of generative UX research for an industrial NLP application for the process industry. Section 5 describes the implementation of the prototype, i.e., semantic search in a complex domain-specific German language of the daily operations. Section 6 reports prototype evaluation with domain users assessing it from two perspectives: (1) accuracy evaluation with information retrieval metrics, (2) user value estimation through the prototype usability evaluation. Section 7 discusses the impact of the generative UX research as a part of developing NLP prototypes of domain-specific applications. Finally, we summarize the findings into a reusable methodology for applying generative research to increase the quality of NLP domain applications.
@@ 2 Related Work
User experience research has become a requirement for practical and effective software development [21, 32]. UX puts perspective onto users, thus defining functionality and look of new products or features and ensuring a need in new products [1]. UX research methods analyze how users interact with a system of interest and explore what users miss or find inconvenient. Quantitative (e.g., questionnaires) and qualitative (e.g., interviews) methods are used to collect and analyze user feedback [43]. Structured interviews with domain-specific probe questions help UX to elicit knowledge about the domain essential for a better understanding of system requirements [28].
@@
The rapid growth of ML applications evolved from automating routine processes (e.g., spam filtering) to user-centered applications (i.e., automatic information extraction based on a specific user request). Although ML applications put users into the loop to ensure the usability of the application [50, 18], most of the ML user-oriented development is product-focused, i.e., the rather later stages engage users to evaluate product usability [47]. UX-driven research embeds user perspective evenly during the development of ML-powered applications, i.e., at the stages of requirement understanding [12], data collection [38], implementation [14], and evaluation [11]. The UX research explores multiple aspects of user interaction and engagement in the development process. UX research increases the effectiveness and communicability of ML applications according to the user’s mental maps [3], creates workflows to enable effective joint creativity process of users and ML models [31], ensures the best user experience from the services [18], incorporates expert knowledge into improving the ML applications [4, 19], and engages users into accuracy evaluation [11, 9].
@@
Although users participate in multiple steps of the development process, a deeper user-centered perspective is required to distill users’ real needs [18]. Most UX studies for ML applications follow a one- or two-step product-centric process [47, 51]: (1) interact with the users to learn about a domain, (2) perform usability evaluation of an ML application. Very little research integrates the domain users into the first step to explore their perspectives for idea generation and concept analysis. Exploring user perspectives would include learning about the domain users; which roles they have in this domain; how they use a currently existing system (if any); and what needs and pain points the users may have. Analysis of this information could evolve the ML development into a user-centric viewpoint.
@@
Generative research, which is often applied in design and software development, aims at learning domain user stories and personal experience through exploratory, generative, and evaluative methods [27]. Generative research targets a deeper understanding of user needs and develops product concepts based on the discovered user goals, values, and explanations of their behavior [5].
In this paper, we explore how a methodology of generative UX research could be applied to the full-cycle prototype development of a domain-specific NLP application, i.e., from concept ideation to prototype evaluation. We investigate this research question through a case study of developing an NLP application to increase effectiveness and efficiency in the daily operations of the process industry. We build our study on the combined methodologies proposed by [27] and [13] and reinforce it with domain understanding through data analysis and user interviews.
@@ 3 Methodology
Generative research encourages user-focused development by getting to know the users better and explores opportunities for an ML/NLP application by learning users’ stories, pain points, and needs [47, 5]. A downside of the existing methodologies is the lack of coverage of the full-cycle development. They either describe how to learn about user needs and experiences [5], create a product concept based on the learned information from users and ask them to evaluate this concept [27], or employ users to evaluate a prototype before developing a product for the production-ready state [13]. Moreover, these methodologies do not consider an aspect of data-driven research that should be taken into account when creating an NLP application.
Figure 1: The proposed methodology of generative UX research for end-to-end prototyping of domain NLP applications. The methodology consists of three stages: exploration (data exploration and generative research), prototyping (development of the first MVP (minimal viable product), and evaluation (assessment of accuracy and dynamics in user experience).
We propose combining the above-mentioned methodologies for full-cycle prototype development via generative research that combines UX+ML/NLP to (1) explore ideas for an NLP application based on the stories of domain users, (2) collect feedback on a concept of a proposed NLP solution to their identified needs or pain points, (3) conduct a user-driven evaluation of a prototype from perspectives of user experience with a new NLP solution and accuracy estimation. Figure 1 depicts the proposed methodology in the form of a timeline that we followed to develop and evaluate a domain-specific NLP application. The methodology consists of three steps:
@@
    exploration, i.e., design ideas for a prototype through (1) exploratory data analysis, (2) generative UX research via contextual inquiries, (3) moderated concept evaluation,
@@
    prototyping to develop a fully working prototype with back- and frontend,
@@
    unmoderated prototype evaluation, i.e., evaluation of dynamics in user experience and accuracy assessment
The methodology combines both data- and user-driven approaches that correspondingly come from ML/NLP and UX research. With a team consisting of UX designers, data scientists, and developers, we ensure diverse and effective collaboration for each stage in our pipeline [52]. Discussing prototype ideas requires a balanced view on project aspects such as technical capabilities of NLP approaches and models, possible limitations (e.g., data or computation power limitations), and final value to domain users [51, 12]. The ultimate goal of the methodology is to increase user involvement and interest in an industrial application [37, 9].
4 Exploration: Generative research
Exploring ideas for an NLP prototype aims at collecting the necessary information to understand a domain in general, get familiar with the domain data, and learn about the user’s needs. We tailor a methodology for concept creation via generative research proposed by [27] to a more specific use-case of concept ideation for a domain-specific industry application. Our generative research consists of three steps:
@@ 4 Exploration: Generative research
    exploratory data analysis: learn about the domain system and data, create hypotheses that we could probe during user interviews.
@@
    idea generation through user interviews: (1) conduct user interviews in a focus group to learn about their daily routines and systematize user interviews to identify patterns, and (2) generate a concept for an NLP application or feature.
@@
    concept evaluation: perform moderated concept evaluation of a click-dummy, i.e., conduct a second round of user interviews to test the usability and usefulness of the proposed concept.
4.1 Domain understanding and exploratory data analysis
A goal of domain understanding and exploratory data analysis is to learn about the area for the prototype and investigate data caveats that can either support or limit various stages of development. Our steps for domain exploration include both qualitative analysis, i.e., examining text data and identifying major roles of the domain users, and quantitative analysis, i.e., statistical analysis of the text records and composition of domain vocabulary.
@@ 4.1 Domain understanding and exploratory data analysis
We implement an NLP feature for an existing system of shiftbooks that is used for logging information about daily operations in the process industry, e.g., chemical or pharmaceutical production. Integrating a feature into an existing system involves conceptualizing difficulties, such as adhering to certain workflows and users’ mind maps [51]. We conduct a case study with one plant in the chemical domain that operates in the German language and utilize their database for data analysis and prototype development.
@@
Figure 2: A simplified example of a record in a logging system for the daily operations in the process industry. A record is semi-structured, i.e., it has structured attributes of the records and unstructured text descriptions. The domain language contains professional terms, abbreviations, shortenings, chemical formulas, various IDs (e.g., order IDs), and inventory codes of the plant machinery and equipment. The example is in English for better readability.
Figure 3: Quantitative analysis of the domain text records of a logging system. Left: A histogram of the length of text records shows that most of the records are very short, i.e., smaller than 200 chars. Right: The domain vocabulary includes not only words but also a significant number of codes of functional locations (i.e., machinery) that act like nouns. Additionally, the vocabulary contains a lot of digit-containing terms that refer to IDs or numerical values, e.g., measurements.
We observed two significant signs of a domain language. First, records contain quite a diverse writing style. While some records contained complete sentences, some records looked more like handwritten notes about statuses and a list of tasks to address later. Second, the text contains numeric data, abbreviations, product names, special symbols, terminology related to chemistry, and encoded terms that act like words. For example, Figure 2 shows that the terms “P5002” and “R105.12” act as nouns that correspondingly refer to a pump and a reactor. But outside of this domain, it is impossible to know the meaning of these terms. Moreover, the abbreviation TSR, which stands for technical safety review, and “maint”, which stands for “maintenance,” illustrate the short word forms used in this domain.
@@
We performed quantitative data analysis to explore the findings from the qualitative analysis further and compare a general domain with the production-specific domain terminology. The analysis included the following tasks: (a) evaluate the length of the text records, (b) inspect text composition, and (c) explore the difference between the domain language and the general domain.
@@
Figure 3 shows a histogram of the text lengths, where the majority of the records do not exceed 200 chars, i.e., the size of Twitter posts. A pie chart depicts the proportion of the types of terms based on the overall number of terms in a dataset for analysis. We observed a significant impact of the numeric values (10%) and encoded terms of the local domain (16%), i.e., abbreviations for the functional locations used in the production. Such terms require local dictionaries to enable decoding the terms for out-of-domain understanding.
To explore the prevalence of the domain terminology, we visualized terms from the local domain and general use in a scatter plot. For the local domain, we extracted 1000 most frequent nouns from 20 000 randomly selected text records. For the general domain, we used a dump of German Wikipedia  [^1] https://github.com/GermanT5/wikipedia2corpus and also extracted 1000 most frequent nouns from the randomly selected 20 000 sentences. We focused on nouns as the most content-bearing terms. Figure 4 depicts t-SNE projection of all extracted terms vectorized with fastText word embeddings [36] and shows that domain and general terms lean towards different parts of the plot. Moreover, on the right plot, we see that some of the frequent terms do not have any overlap with the general domain and are unique. Indeed, some domain terms do not occur anywhere in Wikipedia (e.g., “Trennschichtsonde,” “Einstiegsbereit,” and “Laderohr”).
@@
Our qualitative and quantitative analyses lead to the following findings: (1) Text quality is quite uneven compared to the regular text, i.e., short texts of Twitter size often have the format of handwritten notes with incomplete sentences. (2) The text data contains many domain terms, i.e., specific terminology for the process and chemical industry, domain abbreviations, product names, and codes of functional locations that act like nouns. We hypothesize that text records with such uneven structure and writing style most likely have a short lifetime and are mainly used to keep a track of the current processes and production state, rather than being used as a source of knowledge collected over the long term.
4.2 Idea generation through user interviews
The objective of generative research is to learn how domain users use the system on a daily basis and elicit ideas for an NLP-based solution to improve the effectiveness of their existing processes. We perform contextual inquiries, i.e., one-to-one interviews with a group of representative domain users in their own environment [7]. We organized a week-long workshop at a plant to ensure an immersed environment for the research team [45, 7].

### v4

@@ Abstract
User experience (UX) is a part of human-computer interaction (HCI) research and focuses on increasing intuitiveness, transparency, simplicity, and trust for the system users. Most UX research for machine learning (ML) or natural language processing (NLP) focuses on a data-driven methodology. It engages domain users mainly for usability evaluation. Moreover, more typical UX methods tailor the systems towards user usability, unlike learning about the user needs first. This paper proposes a new methodology for integrating generative UX research into developing domain NLP applications. Generative UX research employs domain users at the initial stages of prototype development, i.e., ideation and concept evaluation, and the last stage for evaluating system usefulness and user utility. The methodology emerged from and is evaluated on a case study about the full-cycle prototype development of a domain-specific semantic search for daily operations in the process industry. A key finding of our case study is that involving domain experts increases their interest and trust in the final NLP application. The combined UX+NLP research of the proposed method efficiently considers data- and user-driven opportunities and constraints, which can be crucial for developing NLP applications.
@@
generative research, natural language processing, human-computer interaction, user experience, domain-specific applications
@@ 1 Introduction
Often, the data-driven approach falls short in exploring the needs of the domain users [58]. On the one hand, domain users are often integrated into development at the late test phase to evaluate the usability of ML/NLP applications [13]. Unlike user-driven software development, the development of NLP applications depends mainly on data availability or experimenting with machine learning (ML)/NLP trends and thus is a major driver of application development. On the other hand, the user-driven development of a domain-specific ML/NLP application in medicine showed that close collaboration with the domain users in the earlier stages increases the effectiveness of the final product [57]. Therefore, integrating user experience (UX) and human-computer interaction (HCI) research into ML/NLP research addresses users’ needs, fuses their expertise, and increases intuitiveness, transparency, simplicity, and trust for the system users [8, 44].
This paper reports on developing a real-world domain NLP application for improving effectiveness and efficiency in daily plant operations in the process industry. A major project goal was to achieve high user acceptance for the final application, which required involving users from the initial design phase. That resulted in the proposed methodology for integrating generative research into a development cycle of domain NLP applications by combining data- and user-driven concept creation, prototype development, and prototype evaluation. Widely applied in design and software development, generative UX research is a methodology that aims to learn about domain users and their stories [29, 4]. To increase the credibility and utility of the final application, generative research engages the users in the early stage of development.
@@
The proposed methodology combines data- and user-driven approaches to full-cycle UX-driven prototype development for NLP applications. Prototyping or developing the first minimal viable products (MVPs) most benefits from the methodology. The proposed methodology consists of the following phases. First, the methodology describes generative UX research steps to explore opportunities for ideas and application concepts. This phase relies on exploratory data analysis and contextual inquiries, i.e., user interviews and on-site observations, as a source of idea generation. Furthermore, the domain users validate a proposed concept and help define the most important features for prototype implementation. Second, the prototyping addresses the most important identified research challenges crucial for usability assessment. Lastly, the prototype evaluation measures user utility via contextual inquiries before and after interacting with the prototype. We provide a project timeline with UX- and NLP-specific stages and tasks as a template for developing domain NLP applications.
@@
Our team comprises UX researchers, data scientists, software developers, and a project leader because such a team composition facilitates a deeper understanding of the capabilities and limitations of the domain NLP application [19]. Therefore, the paper may have specific interest for the individual practitioners with the abovementioned roles or similar teams. Moreover, practitioners of user-driven research, e.g., customer research departments and quality assurance officers, could benefit from the proposed methodology and reported results.
@@
The paper is structured as follows. Section 2 briefly reviews previous efforts and UX methodologies in creating ML applications. Section 3 describes a methodology of generative UX research, in which we propose to optimize user experience in the usually data-driven development of domain NLP applications. Our case study’s methodology covers a full-cycle NLP prototype development from generative research for concept ideation to user-driven prototype evaluation. Section 4 reports a case study about implementing the proposed methodology to prototype an industrial NLP application for the process industry. Section 5 discusses the impact of the generative UX research on developing our NLP prototype and summarizes the findings into a reusable methodology for applying generative research to increase the quality of NLP domain applications.
@@ 2 Related Work
User experience research has become a requirement for practical and effective software development [22, 37]. UX puts perspective onto users, thus defining functionality and look of new products or features and ensuring a need in new products [26]. UX research methods analyze how users interact with a system of interest and explore what users miss or find inconvenient. Quantitative (e.g., questionnaires) and qualitative (e.g., interviews) methods are used to collect and analyze user feedback [49]. Structured interviews with domain-specific probe questions help UX elicit knowledge about the domain essential for better understanding system requirements [32].
Domain-driven design (DDD) is a widely used methodology of software development that aims to reflect real world domains through the architecture and interaction within and with the system [43, 36, 53, 30, 31]. UX focuses on learning about domains and domain-specific tasks to identify actors, entities, and processes and formalize them as a preliminary step of software engineering.
@@
The rapid growth of ML applications evolved from automating routine processes (e.g., spam filtering) to user-centered applications (i.e., automatic information extraction based on a specific user request). Although ML applications put users into the loop to ensure the usability of the application [57, 19], most of the ML user-oriented development is product-focused, i.e., the rather later stages engage users to evaluate product usability [54]. UX-driven research embeds user perspective evenly during the development of ML-powered applications, i.e., at the stages of requirement understanding [12], data collection [45], implementation [14], and evaluation [11]. The UX research explores multiple user interaction and engagement aspects in the development process. UX research increases the effectiveness and communicability of ML applications according to the user’s mental maps [2], creates workflows to enable effective joint creativity process of users and ML models [35], ensures the best user experience from the services [19], incorporates expert knowledge into improving the ML applications [3, 20], and engages users into accuracy evaluation [11, 8].
@@
Although users participate in multiple steps of the development process, a deeper user-centered perspective is required to distill users’ real needs [19]. Most UX studies for ML applications follow a one- or two-step product-centered process [54, 58]: (1) interact with the users to learn about a domain, (2) perform usability evaluation of an ML application. Very little research integrates the domain users into the first step to explore their perspectives for idea generation and concept analysis. Exploring user perspectives would include learning about the domain users, their roles in this domain, how they use a currently existing system (if any), and what needs and pain points the users may have. Analysis of this information could evolve the ML development into a user-centered viewpoint.
@@
Generative research, often applied in design and software development, aims to learn domain user stories and personal experiences through exploratory, generative, and evaluative methods [29]. Generative research targets a deeper understanding of user needs and develops product concepts based on the discovered user goals, utility, and underlying reasons for their actions [4]. Generative research is conceptually close to the practical design thinking methodology often applied in industry [9, 38, 18]. Although the ideation process in both methodologies is user-centered, generative research focuses more on exploring domain users in their environment and deriving prototype ideas based on immersive domain learning.
@@ 3 Methodology
In this section, we propose a methodology of generative UX research for optimizing a data-driven domain-specific NLP application toward higher user satisfaction. Generative research encourages user-focused development by getting to know the users better and explores opportunities for an ML/NLP application by learning users’ stories, pain points, and needs [54, 4]. Although generative research has already been employed in design and software development, it lacks coverage of full-cycle development of ML/NLP applications. For example, [4] focused on how to learn about user needs and experiences. [29] explored how to create a product concept based on the learned information from users and employed the users to evaluate this concept. Similarly, design thinking proposes a user-centered process of iterative ideating and prototyping [9, 38, 18]. [13] employed users to evaluate a prototype before developing a product for the production-ready state. Moreover, these methodologies do not consider an aspect of data-driven research that should be considered when creating an NLP application. To enable the full-cycle prototype development, the proposed methodology combines these previous efforts on generative research, design thinking, and domain understanding through exploratory data analysis required in ML/NLP applications.
Figure 1: Project timeline for the proposed methodology of generative UX research for prototyping domain NLP applications used in our case study. The methodology consists of three stages: exploration (data exploration and generative research), prototyping (development of the prototype and evaluation (assessment of accuracy and user utility). The methodology combines UX and NLP tasks, encouraging the teams to collaborate closely, especially during the generative research phase. The timeline reflects the reported case study’s duration, but each phase’s duration may differ when applied to other projects.
The methodology for full-cycle prototype development through generative research combines UX+ML/NLP to (1) explore ideas for an NLP application based on the stories of domain users, (2) collect feedback on a concept of a proposed NLP solution to their identified needs or pain points, (3) conduct a user-driven evaluation of a prototype from perspectives of user experience with a new NLP solution and accuracy estimation. Figure 1 depicts the proposed methodology as a timeline we followed to develop and evaluate a domain-specific NLP application. The methodology consists of three steps:
@@
    Exploration, i.e., design ideas for a prototype through exploratory data analysis, interviews, and observation of the domain users in their environment.
@@
    Prototyping to develop a fully working prototype with back- and frontend.
@@
    Prototype evaluation, i.e., evaluation of the change in user experience and prototype accuracy assessment.
Exploring ideas for an NLP prototype through generative UX research aims to collect the necessary information to understand a domain in general, get familiar with the domain data, and learn about the user’s needs. We tailor a methodology for concept creation via generative research proposed by [29] to a more specific use-case of concept ideation for a domain-specific industry application. A proposed generative research for NLP consists of three steps:
@@
    Exploratory data analysis: Learn about the domain system and data and create hypotheses we could probe during user interviews.
@@
    Contextual inquiries: (1) Conduct user interviews in a focus group to learn about their daily routines and systematize user interviews to identify patterns, and (2) Generate ideas to propose a concept for an NLP application or feature.
@@
    Concept evaluation: Perform moderated concept evaluation of a click-dummy, i.e., enable supervised interaction with a “click-dummy” prototype to test the usability and usefulness of the proposed concept.
The user-driven prototype evaluation aims to assess prototype accuracy and how user utility changes after interacting with the prototype. Unlike concept evaluation at the stage of generative UX research, the prototype is evaluated in an open environment. The combination of the two perspectives (i.e., accuracy and user utility) determines the success level of the prototype in the first step toward a system for shared domain knowledge. The evaluation consists of three steps [13]:
  1. 1.
  2. 2.
    Experimental: Provide the prototype to interact with it through provided tasks and openly throughout daily tasks as they feel fit;
  3. 3.
    Reflection: Reuse the same questionnaire and interview scripts to identify how the user experience changes and learn if the system works as expected or surprises them and provides insights.
The methodology combines data- and user-driven approaches that correspondingly originate from ML/NLP and UX research. A team of UX designers, data scientists, and developers ensures diverse and effective collaboration for each stage in the proposed pipeline [59]. Discussing prototype ideas requires a balanced view of project aspects such as technical capabilities of NLP approaches and models, possible limitations (e.g., data or computation power limitations), and benefit to domain users [58, 12]. The ultimate goal of the methodology is to increase user involvement and interest in an industrial application and increase system usefulness [44, 8].
4 Case study
This case study describes implementing a domain-specific NLP application into a digital tool for capturing shift logs and managing plant operations in the chemical process industry. Shift logs usually contain problems and solutions arising during production, performed tasks, instructions from supervisors, reports about production goals, changes on production lines, and much more information about daily operations. As the system installed for a study participant plant had been used for several years, a large collection of text logs was available as a data source for NLP applications.
We consulted a user subgroup that closely interacts with the system daily to ideate how an NLP feature or application would look and evaluate which solution would be most effective for the domain system users. The main challenge was that the way the users intended to use the NLP application within the overall scope of tasks did not coincide with the data shape or form. Therefore, the goal of the case study was to devise a pilot idea based on the user perspective and derive an application that will assist domain users’ daily tasks based on the data at hand. For this, we employed and implemented our proposed methodology of generative UX research for full-cycle prototyping domain NLP applications.
4.1 Exploration
4.1.1 Domain understanding and exploratory data analysis
A goal of domain understanding and exploratory data analysis is to learn about the area for the prototype and investigate data caveats that can either support or limit various stages of development. Our domain exploration steps included qualitative analysis, i.e., examining text data and identifying major roles of the domain users, and quantitative analysis, i.e., statistical analysis of the text records and composition of domain vocabulary.
@@ 4.1.1 Domain understanding and exploratory data analysis
We implemented an NLP feature for an existing system of electronic shift books, which is used for logging information about daily operations in the process industry, e.g., chemical or pharmaceutical production. Integrating a feature into an existing system involves conceptualizing difficulties, such as adhering to certain workflows and users’ mind maps [58]. We conducted a case study with one plant in the chemical domain that operates in the German language. A database with the plant’s daily operations logs is used for data analysis and prototype development.
@@
Figure 2: A simplified example of a record in a logging system for the daily operations in the process industry. A record is semi-structured, i.e., it has structured attributes of the records and unstructured text descriptions. The domain language contains professional terms, abbreviations, shortenings, chemical formulas, various IDs (e.g., order IDs), and plant machinery and equipment inventory codes. The example is in English for better readability.
Figure 3: Quantitative analysis of the domain text records of a logging system. Left: A histogram of the length of text records shows that most records are very short, i.e., smaller than 200 chars. Right: The domain vocabulary includes words and many codes of functional locations (i.e., machinery) that act like nouns. Additionally, the vocabulary contains a lot of digit-containing terms that refer to IDs or numerical values, e.g., measurements.
We observed two significant signs of a domain language. First, records contain quite a diverse writing style. While some records contained complete sentences, some looked like handwritten notes about statuses and a list of tasks to address later. Second, the text contains numeric data, abbreviations, product names, special symbols, terminology related to chemistry, and encoded terms that act like words. For example, Figure 2 shows that the terms “P5002” and “R105.12” act as nouns that correspondingly refer to a pump and a reactor. But outside of this domain, knowing the meaning of these terms is impossible. Moreover, the abbreviation TSR, which stands for technical safety review, and “maint,” which stands for “maintenance,” illustrate the short word forms used in this domain.
@@
We performed quantitative data analysis to explore the qualitative analysis findings further and compare a general domain with the production-specific domain terminology. The analysis included the following tasks: (a) evaluate the length of the text records, (b) inspect text composition, and (c) explore the difference between the domain language and the general domain.
@@
Figure 3 shows a histogram of the text lengths, where most records do not exceed 200 chars, i.e., the size of Twitter posts. A pie chart depicts the proportion of the types of terms based on the overall number of terms in a dataset for analysis. We observed a significant impact of the numeric values (10%) and encoded terms of the local domain (16%), i.e., abbreviations for the functional locations used in the production. Such terms require local dictionaries to enable decoding the terms for out-of-domain understanding.
To explore the prevalence of the domain terminology, we visualized terms from the local domain and general use in a scatter plot. For the local domain, we extracted 1000 most frequent nouns from 20,000 randomly selected text records. For the general domain, we used a dump of German Wikipedia  [^1] https://github.com/GermanT5/wikipedia2corpus and extracted the 1000 most frequent nouns from the randomly selected 20,000 sentences. We focused on nouns as the most content-bearing terms. Figure 4 depicts a t-SNE projection of all extracted terms vectorized with fastText word embeddings [42] and shows that domain and general terms lean towards different parts of the plot. Moreover, on the right plot, we see that some frequent terms are unique and do not overlap with the general domain. For example, the terms “Trennschichtsonde,” “Einstiegsbereit,” and “Laderohr” do not occur in Wikipedia.
@@
Our qualitative and quantitative analyses led to the following findings: (1) Text quality was quite uneven compared to the regular text, i.e., short texts of Twitter size often have the format of handwritten notes with incomplete sentences. (2) The text data contained many domain terms, i.e., specific terminology for the process and chemical industry, domain abbreviations, product names, and codes of functional locations that act like nouns. We hypothesized that text records with such uneven structure and writing style most likely have a short lifetime and are mainly used to keep track of the current processes and production state rather than as a source of knowledge collected over the long term.
4.1.2 Idea generation through user interviews
The objective of generative research is to learn how domain users use the system daily and elicit ideas for an NLP-based solution to improve the effectiveness of their existing processes. We performed contextual inquiries, i.e., one-to-one interviews with representative domain users in their environment [6]. We organized a week-long workshop at a plant to ensure an immersed environment for the research team [51, 6].